### v1

# Classification of locality preserving symmetries on spin chains

Abstract
We consider the action of a finite group $G$ by locality preserving automorphisms (quantum cellular automata) on quantum spin chains. We refer to such group actions as “symmetries”. The natural notion of equivalence for such symmetries is stable equivalence, which allows for stacking with factorized group actions. Stacking also endows the set of equivalence classes with a group structure. We prove that the anomaly of such symmetries provides an isomorphism between the group of stable equivalence classes of symmetries with the cohomology group $H^{3}(G,U(1))$, consistent with previous conjectures. This amounts to a complete classification of locality preserving symmetries on spin chains. We further show that a locality preserving symmetry is stably equivalent to one that can be presented by finite depth quantum circuits with covariant gates if and only if the slant product of its anomaly is trivial in $H^{2}(G,U(1)[G])$.
Appendix A Introduction
Dynamics in many-body quantum physics is typically generated by a local Hamiltonian, and therefore, due to Lieb-Robinson bounds [1], it preserves locality. Such locally generated evolutions may be thought of as topologically trivial locality preserving automorphisms of the observable algebra. Indeed, they are contracted to the identity by reducing the evolution time. In contrast, the shift on a spin chain is an example of a topologically non-trivial locality preserving automorphism [2].
Non-trivial locality preserving automorphisms appear in various guises in the study of topological phases of strongly interacting quantum matter, and often serve to characterize and even classify the phases under investigation [3]. Examples include the appearance at stroboscopic times of the shift on the boundary of many-body localized Floquet insulators [4, 5, 6], the exotic symmetries appearing on the boundaries of topological matter [7, 8], and the equivariant automorphisms that can entangle symmetry protected trivial (SPT) phases [9, 6]. These connections have motivated a growing body of work that aims to understand the topological phases of locality preserving automorphisms, possibly in the presence of symmetry [2, 10, 11, 12, 13, 14, 15, 16].
In this paper we study representations of finite groups $G$ by locality preserving automorphisms on spin chains. Such locality preserving symmetries arise at the boundaries of two-dimensional SPTs whose bulk invariant manifests itself as an $H^{3}(G,U(1))$-valued anomaly of the boundary symmetry [17]. We prove the folk knowledge that this anomaly classifies locality preserving symmetries on spin chains up to stable equivalence. That is, up to conjugation by finite depth quantum circuits and stacking with factorized group actions.
Our proof proceeds by first asking which locality preserving symmetries on spin chains admit restrictions to right half-lines that are
  1. (1)
    themselves locality preserving symmetries,
  2. (2)
    covariant with respect to the full symmetry.
The anomaly is an obstruction to (1). We introduce in addition a new obstruction to (2), called the obstruction to covariant right restrictions, which takes values in twisted group cohomology $H^{2}(G,U(1)[G])$. We then show that a symmetry with trivial anomaly also has trivial obstruction to covariant right restrictions, and that the existence of right restrictions that satisfy (1) and (2) simultaneously implies that symmetries with trivial anomaly can be decoupled. The solution of the classification problem then follows from the fact that the anomaly and the obstruction to covariant right restrictions are constant on stable equivalence classes of symmetries.
As a corollary we show that the obstruction to covariant right restrictions is given by the inverse of the slant product of the anomaly. Having a good handle on this quantity is significant because it plays an important role in characterizing the anyon content of the gauged bulk SPT corresponding to the boundary symmetry under consideration [18, 19]. Note in particular that any symmetry which admits covariant right restrictions can be presented by finite depth quantum circuits with covariant gates, a highly non-trivial property.
The paper is structured as follows. In Section B we introduce locality preserving symmetries on spin chains and stable equivalence between them, and state our main Theorem. We define the anomaly in Section C and state its basic properties. In Section D we construct for each element of $H^{3}(G,U(1))$ an explicit symmetry with that element as its anomaly. In Section E we define the obstruction to covariant right restrictions and show that it vanishes for symmetries with trivial anomaly. This fact is then used in Section F to prove the main Theorem. Appendix G collects basic definitions of group cohomology. Basic properties of the anomaly and of the obstruction to covariant right restrictions are proved in Appendices H and I respectively. In Appendix J we show that the obstruction to covariant right restrictions is given by the slant product of the anomaly. Finally, Appendix K presents an example which shows that stable equivalence is needed in order for the classification by the anomaly to hold.
Note : During the preparation of this manuscript, the preprint [20] appeared, in which similar results are obtained. In particular, the disentangler $\mathcal{W}$ constructed in [20] yields a proof of our Proposition F.1.
Acknowledgments : W.D.R. and B.O.C. were supported by the FWO (Flemish Research Fund) grant G098919N, the FWO-FNRS EOS research project G0H1122N EOS 40007526 CHEQS, the KULeuven Runners-up grant iBOF DOA/20/011, and the internal KULeuven grant C14/21/086.
Appendix B Setup and main result
B.1 Spin chains, quantum cellular automata, and finite depth quantum circuits
A spin chain is an assignment of a full matrix algebra ${\mathcal{A}}_{j}\simeq\operatorname{\mathrm{End}}({\mathbb{C}}^{d_{j}})$ to each site $j\in{\mathbb{Z}}$ of the chain with uniformly bounded dimensions $d_{j}$. We write ${\mathcal{A}}$ for the corresponding algebra of quasi-local observables, and for any $\Gamma\subset{\mathbb{Z}}$ we write ${\mathcal{A}}_{\Gamma}$ for the $C^{*}$-algebra of quasi-local observables supported in $\Gamma$. We will write ${\mathcal{A}}_{\geq j}$ for ${\mathcal{A}}_{[j,\infty)}$ and ${\mathcal{A}}_{<j}$ for ${\mathcal{A}}_{(-\infty,j-1]}$. We will usually refer to the quasi-local algebra ${\mathcal{A}}$ itself as the ‘spin chain’, it being understood that there is a fixed preferred assignment of on-site algebras $j\mapsto{\mathcal{A}}_{j}\subset{\mathcal{A}}$ to sites of ${\mathbb{Z}}$.
For any $\Gamma\subset{\mathbb{Z}}$ we write $\Gamma^{(r)}:=\{j\in{\mathbb{Z}}\,:\,\operatorname{\mathrm{dist}}(j,\Gamma)\leq r\}$ for the ‘$r$-fattening’ of $\Gamma$.
A quantum cellular automaton (QCA) on a spin chain ${\mathcal{A}}$ is a *-automorphism $\alpha:{\mathcal{A}}\rightarrow{\mathcal{A}}$ for which there exists $r\geq 0$ such that $\alpha(x)\in{\mathcal{A}}_{\mathrm{supp}\,(x)^{(r)}}$ for all $x\in{\mathcal{A}}$.
The range of a QCA is the smallest $r$ for which this holds. The inverse of a QCA of range $r$ is also a QCA of range $r$ ([10, Lemma 3.1]), a fact which we will use without comment throughout the paper. The quantum cellular automata on ${\mathcal{A}}$ form a subgroup of $\operatorname{\mathsf{Aut}}({\mathcal{A}})$ which we denote by $\operatorname{\mathsf{QCA}}({\mathcal{A}})$.
Let $\{I_{a}\}_{a\in{\mathbb{Z}}}$ be a partition of ${\mathbb{Z}}$ into intervals $I_{a}\subset{\mathbb{Z}}$ of bounded size. Suppose we have for each $a\in{\mathbb{Z}}$ a unitary $U_{a}\in{\mathcal{A}}_{I_{a}}$, then we can define a QCA $\beta$ which is given by conjugation with the (formal) unitary $\bigotimes_{a\in{\mathbb{Z}}}U_{a}$. Any QCA of this form is called a block partitioned QCA. The intervals $I_{a}$ are called the blocks of the block partitioned QCA, and $\left|I_{a}\right|$ is the size of block $I_{a}$. The unitaries $U_{a}$ are called gates. The composition of $n$ block partitioned QCAs is called a depth $n$ quantum circuit, or simply a finite depth quantum circuit (FDQC).
B.2 Locality preserving symmetries
Let $G$ be a finite group which will be fixed throughout the paper. We write $\bar{g}=g^{-1}$ for the inverse of any group element $g\in G$. A locality preserving symmetry on ${\mathcal{A}}$ is a group homomorphism $\alpha:G\rightarrow\operatorname{\mathsf{QCA}}({\mathcal{A}})$. That is, for each $g\in G$ we have a quantum cellular automaton $\alpha^{(g)}$ such that $\alpha^{(1)}=\mathrm{id}$ and $\alpha^{(g)}\circ\alpha^{(h)}=\alpha^{(gh)}$ for all $g,h\in G$. The range of a locality preserving symmetry is the largest range of its component QCAs. We say a locality preserving symmetry $\alpha$ is decoupled if there is a partition $\{I_{a}\}_{a\in{\mathbb{Z}}}$ of ${\mathbb{Z}}$ into intervals $I_{a}$ of bounded size such that $\alpha=\bigotimes_{a\in{\mathbb{Z}}}\,\alpha_{a}$ for symmetries $\alpha_{a}$ supported on $I_{a}$. In particular, any symmetry of range 0 is decoupled.
In the rest of this work we will refer to locality preserving symmetries simply as symmetries. If we want to specify the group $G$ then we speak of $G$-symmetries. We denote the set of all $G$-symmetries by $\operatorname{\mathsf{Sym}}_{G}$.
B.3 Equivalence and stable equivalence
Two symmetries $\alpha$ and $\alpha^{\prime}$ are equivalent if there is a FDQC $\gamma$ such that $\alpha^{\prime(g)}=\gamma^{-1}\circ\alpha^{(g)}\circ\gamma$ holds for all $g\in G$. In that case we write $\alpha^{\prime}\sim_{0}\alpha$.
The stack of two spin chains ${\mathcal{A}}$ and ${\mathcal{B}}$ is the spin chain ${\mathcal{A}}\otimes{\mathcal{B}}$ with on-site algebras $({\mathcal{A}}\otimes{\mathcal{B}})_{j}={\mathcal{A}}_{j}\otimes{\mathcal{B}}_{j}$ for all $j\in{\mathbb{Z}}$. If $\alpha$ and $\beta$ are symmetries on the spin chains ${\mathcal{A}}$ and ${\mathcal{B}}$ respectively, then we can stack them to obtain the symmetry $\alpha\otimes\beta$ on ${\mathcal{A}}\otimes{\mathcal{B}}$ with components $(\alpha\otimes\beta)^{(g)}=\alpha^{(g)}\otimes\beta^{(g)}$ for all $g\in G$.
Two symmetries $\alpha$ and $\alpha^{\prime}$ are stably equivalent, denoted by $\alpha\sim\alpha^{\prime}$, if there exists decoupled symmetries $\beta$ and $\beta^{\prime}$ such that $\alpha\otimes\beta\sim_{0}\alpha^{\prime}\otimes\beta^{\prime}$. Stable equivalence is an equivalence relation on $\operatorname{\mathsf{Sym}}_{G}$, and $(\operatorname{\mathsf{Sym}}_{G}/\sim)$ is an abelian monoid with multiplication induced by stacking. (We will show later that it is in fact a group, i.e. there are inverses.) Note that any decoupled symmetry is stably equivalent to a symmetry of range zero.
B.4 Main result
Theorem B.1.
The monoid $(\operatorname{\mathsf{Sym}}_{G}/\sim)$ is in fact a group. There is a map $\Omega:\operatorname{\mathsf{Sym}}_{G}\rightarrow H^{3}(G,U(1))$ which assigns to each symmetry $\alpha$ its anomaly and which lifts to an isomorphism of groups $(\operatorname{\mathsf{Sym}}_{G}/\sim)\cong H^{3}(G,U(1))$.
In particular, two $G$-symmetries $\alpha$ and $\beta$ are stably equivalent if and only if their anomalies are equal:
$\alpha\sim\beta\,\iff\,\Omega(\alpha)=\Omega(\beta).$
Moreover, for each $[\omega]\in H^{3}(G,U(1))$ there exists a symmetry whose anomaly is $[\omega]$.
This theorem is proven at the end of Section F.
Remark B.2.
In Appendix K we describe a symmetry $\alpha$ with trivial anomaly $\Omega(\alpha)=[1]$ which is nevertheless not equivalent to a decoupled symmetry. This shows that the notion of stable equivalence is indeed necessary for the classification by the anomaly to hold.
Appendix C The anomaly of a locality preserving symmetry
The idea behind the definition of the anomaly presented here goes back to [17]. In order to define the anomaly we first note that the component QCAs of any locality preserving symmetry are finite depth quantum circuits [21].
Lemma C.1.
Let $\alpha:G\rightarrow\operatorname{\mathsf{QCA}}({\mathcal{A}})$ be a symmetry of range $R$ on a spin chain ${\mathcal{A}}$. Then each $\alpha^{(g)}$ can be written as a depth two quantum circuit whose blocks all have size at most $2R$.
Proof.
For each $g\in G$ we have a QCA $\alpha^{(g)}$ on the spin chain ${\mathcal{A}}$. To any such QCA one can assign its ${\mathbb{Q}}$-valued GNVW index $\operatorname{\mathrm{ind}}(\alpha^{(g)})\in{\mathbb{Q}}$, see [2]. Since $G$ is a finite group, $g$ has finite order. i.e. there is an $n$ such that $g^{n}=1$. Since the GNVW index is multiplicative under composition of QCAs and $\operatorname{\mathrm{ind}}(\mathrm{id})=1$, this implies that $\operatorname{\mathrm{ind}}(\alpha^{(g)})^{n}=\operatorname{\mathrm{ind}}(\alpha^{(g^{n})})=\operatorname{\mathrm{ind}}(\alpha^{(1)})=\operatorname{\mathrm{ind}}(\mathrm{id})=1$ and therefore $\alpha^{(g)}$ has trivial GNVW index. The claim now follows from [2, Theorem 9]. ∎
Let $\alpha$ be a symmetry of range $R$. A right restriction $\alpha_{\geq j}$ of $\alpha$ at $j\in{\mathbb{Z}}$ with defect size $L$ is a family of automorphisms $\alpha_{\geq j}^{(g)}$ such that for any $g\in G$
$\alpha_{\geq j}^{(g)}|_{{\mathcal{A}}_{\geq(j+L)}}=\alpha^{(g)}|_{{\mathcal{A}}_{\geq(j+L)}}\,\text{ and }\,\,\alpha_{\geq j}^{(g)}|_{{\mathcal{A}}_{<j-L}}=\mathrm{id}_{{\mathcal{A}}_{<j-L}}.$
It follows immediately from Lemma C.1 that any symmetry of range $R$ admits right restrictions at all sites with defect size $2R$.
Given a right restriction $\alpha_{\geq j}$ of defect size $L$ there are local unitaries $\operatorname{\Phi}_{j}(g,h)\in{\mathcal{A}}_{[j-L,j+L+R]}$, called fusion operators associated to $\alpha_{\geq j}$, such that
$\alpha_{\geq j}^{(g)}\circ\alpha_{\geq j}^{(h)}=\mathrm{Ad}[\operatorname{\Phi}_{j}(g,h)]\circ\alpha_{\geq j}^{(gh)}.$
These unitaries are uniquely determined by this equation up to phase. They capture the failure of $g\mapsto\alpha_{\geq j}^{(g)}$ to be a group homomorphism.
Using associativity to compute $\alpha_{\geq j}^{(f)}\circ\alpha_{\geq j}^{(g)}\circ\alpha_{\geq j}^{(h)}$ in two different ways one obtains
$\mathrm{Ad}\left[\operatorname{\Phi}_{j}(f,g)\operatorname{\Phi}_{j}(fg,h)\right]\circ\alpha_{\geq j}^{(fgh)}=\mathrm{Ad}\left[\alpha_{\geq j}^{(f)}\big{(}\operatorname{\Phi}_{j}(g,h)\big{)}\operatorname{\Phi}_{j}(f,gh)\right]\circ\alpha_{\geq j}^{(fgh)}.$
It follows that there are phases $\omega_{j}(f,g,h)\in U(1)$ such that
$\operatorname{\Phi}_{j}(f,g)\operatorname{\Phi}_{j}(fg,h)=\omega_{j}(f,g,h)\times\alpha_{\geq j}^{(f)}\big{(}\operatorname{\Phi}_{j}(g,h)\big{)}\operatorname{\Phi}_{j}(f,gh)$ (C.1)
for all $f,g,h\in G$.
Proposition C.2.
The map $\omega_{j}:G^{3}\rightarrow U(1)$ is a 3-cocycle,
$1=\frac{\omega_{j}(g,h,k)\omega_{j}(f,gh,k)\omega_{j}(f,g,h)}{\omega_{j}(fg,h,k)\omega_{j}(f,g,hk)}$
and the corresponding group cohomology class $[\omega_{j}]\in H^{3}(G,U(1))$ depends only on the symmetry $\alpha$. We thus obtain a well defined map
$\operatorname{\Omega}:\operatorname{\mathsf{Sym}}_{G}\rightarrow H^{3}(G,U(1))$
which we call the anomaly. If $\operatorname{\Omega}(\alpha)=[1]$ is the identity element of $H^{3}(G,U(1))$ then we say that $\alpha$ has trivial anomaly.
Moreover, for symmetries $\alpha$ and $\beta$ of range $R$ we have
  1. 1.
    If $\alpha$ is decoupled then $\operatorname{\Omega}(\alpha)=[1]$ is the identity element of $H^{3}(G,U(1))$.
  2. 2.
    The anomaly is locally computable: If $\alpha$ and $\beta$ act on the same spin chain ${\mathcal{A}}$ and there is an interval $I$ of length $18R+$ 1 such that $\alpha|_{{\mathcal{A}}_{I}}=\beta|_{{\mathcal{A}}_{I}}$ then $\operatorname{\Omega}(\alpha)=\operatorname{\Omega}(\beta)$.
  3. 3.
    The anomaly is multiplicative under stacking: $\operatorname{\Omega}(\alpha\otimes\beta)=\operatorname{\Omega}(\alpha)\cdot\operatorname{\Omega}(\beta)$.
  4. 4.
    The anomaly is constant on stable equivalence classes: $\alpha\sim\beta\implies\operatorname{\Omega}(\alpha)=\operatorname{\Omega}(\beta)$.
In particular, the anomaly lifts to a homomorphism of monoids $\operatorname{\Omega}:(\operatorname{\mathsf{Sym}}_{G}/\sim)\rightarrow H^{3}(G,U(1))$.
Proof.
Can be found in Appendix H. ∎
Remark C.3.
We could use left restrictions instead of right restrictions to give an alternative anomaly $\operatorname{\Omega}_{L}(\alpha)$. Then one can check that $\operatorname{\Omega}_{L}(\alpha)=\operatorname{\Omega}(\alpha)^{-1}$.
Appendix D Examples
Let $G$ be a finite group and $\omega:G^{3}\rightarrow U(1)$ a 3-cocycle. We construct a symmetry $\alpha$ with anomaly $[\omega]\in H^{3}(G,U(1))$.
Consider the spin chain with local degrees of freedom ${\mathcal{A}}_{x}\simeq\operatorname{\mathrm{End}}\big{(}{\mathbb{C}}^{\left|G\right|}\big{)}$. Define unitaries $U_{j,j+1}^{(g)}\in{\mathcal{A}}_{\{j,j+1\}}$ by
$U_{j,j+1}^{(g)}|g_{j},g_{j+1}\rangle=\omega(g,g_{j+1},\bar{g}_{j+1}g_{j})|g_{j},g_{j+1}\rangle.$
Note that the $U_{j,j+1}^{(g)}$ commute with each other for all $j\in{\mathbb{Z}}$ and for all $g\in G$.
Define $\alpha^{(g)}$ as the composition $\alpha_{3}^{(g)}\circ\alpha_{2}^{(g)}\circ\alpha_{1}^{(g)}$ of three block partitioned QCAs. The blocks of $\alpha_{1}^{(g)}$ are neighbouring pairs of sites $\{2a,2a+1\}$ and the corresponding gates are $U^{(g)}_{2a,2a+1}$. Similarly, $\alpha_{2}^{(g)}$ has blocks $\{2a-1,2a\}$ and corresponding gates $U_{2a-1,2a}^{(g)}$. Finally, the block partitioned QCA $\alpha_{3}^{(g)}$ has the singletons $\{a\}$ as blocks with the left action $L^{(g)}|h\rangle=|gh\rangle$ as gates. See Figure 1.
{quantikz}
[…]
Figure 1: The FDQC defining $\alpha^{(g)}$.
Let $I=[a,b]\subset{\mathbb{Z}}$ be a finite interval and let $\alpha_{I}^{(g)}$ be the FDQC obtained from $\alpha^{(g)}$ by only retaining the gates that are supported on $I$, see Figure 2 The product of the finite number of gates of $\alpha_{I}^{(g)}$ then defines a unitary $U_{I}^{(g)}$ so that $\alpha_{I}^{(g)}=\mathrm{Ad}[U_{I}^{(g)}]$. Note that $\alpha^{(g)}=\lim_{a\uparrow\infty}\alpha_{[-a,a]}^{(g)}$ in the strong topology.
{quantikz}
[…]
Figure 2: The FDQC (in red) defining $\alpha_{I}^{(g)}$ for $I=[a,b]$.
Lemma D.1.
If $I=[a,b]$ then
$U_{I}^{(g)}U_{I}^{(h)}U_{I}^{(gh)*}=\operatorname{\Phi}_{a}(g,h)\times\operatorname{\Phi}_{b}(g,h)^{*}$
with
$\operatorname{\Phi}_{j}(g,h)=\sum_{g_{j}\in G}\omega(g,h,\bar{h}\bar{g}g_{j})\,|g_{j}\rangle\langle g_{j}|.$
Proof.
Let us act on a product state $|g_{a}\,g_{a+1}\cdots g_{b}\rangle$. Let $f_{j}=\bar{g}_{j}g_{j-1}$ for all $j\in\{a+1,\cdots,b\}$. Then
$U_{I}^{(h)}|g_{a}\,g_{a+1}\cdots g_{b}\rangle=\prod_{j=a+1}^{b}\omega(h,g_{j},f_{j})\times|(hg_{a})\,(hg_{a+1})\cdots(hg_{b})\rangle$
so
$U_{I}^{(g)}U_{I}^{(h)}|g_{a}\,\cdots g_{b}\rangle=\prod_{j=a+1}^{b}\omega(h,g_{j},f_{j})\omega(g,hg_{j},f_{j})\times|(ghg_{a})\,\cdots(ghg_{b})\rangle$
and
$U_{I}^{(g)}U_{I}^{(h)}U_{I}^{(gh)*}|(ghg_{a})\cdots(ghg_{b})\rangle$ $=\prod_{j=a+1}^{b}\frac{\omega(h,g_{j},f_{j})\omega(g,hg_{j},f_{j})}{\omega(gh,g_{j},f_{j})}\times|(ghg_{a})\,\cdots(ghg_{b})\rangle$
$=\prod_{j=a+1}^{b}\frac{\omega(g,h,g_{j-1})}{\omega(g,h,g_{j})}\times|(ghg_{a})\,\cdots(ghg_{b})\rangle$
$=\frac{\omega(g,h,g_{a})}{\omega(g,h,g_{b})}\times|(ghg_{a})\,\cdots(ghg_{b})\rangle$
$=\operatorname{\Phi}_{a}(g,h)\operatorname{\Phi}_{b}(g,h)^{*}\times|(ghg_{a})\,\cdots(ghg_{b})\rangle$
where we recalled that $f_{j}=\bar{g}_{j}g_{j-1}$ and we used the cocycle relation (Proposition C.2) in the second step. ∎
It follows immediately from this lemma that $\alpha^{(g)}\alpha^{(h)}=\alpha^{(gh)}$ for all $g,h\in G$, so that $g\mapsto\alpha^{(g)}$ is indeed a group homomorphism.
Let $\alpha_{\geq}^{(g)}:=\lim_{b\uparrow\infty}\alpha_{[0,b]}^{(g)}$. Then $\alpha_{\geq}$ is a right restriction of $\alpha$ and
$\alpha_{\geq}^{(g)}\alpha_{\geq}^{(h)}=\mathrm{Ad}[\operatorname{\Phi}_{0}(g,h)]\alpha_{\geq}^{(gh)}.$
We can now compute the anomaly of $\alpha$. Note that the fusion operators $\operatorname{\Phi}_{j}(g,h)$ commute with all the gates $U_{j,j+1}(f)$, indeed, all these operators are diagonal in the ‘group basis’. We therefore find
$\alpha_{\geq}^{(f)}\big{(}\operatorname{\Phi}_{0}(g,h)\big{)}=\sum_{g_{0}\in G}\,\omega(g,h,\bar{h}\bar{g}g_{0})\,|fg_{0}\rangle\langle fg_{0}|=\sum_{g_{0}\in G}\,\omega(g,h,\bar{h}\bar{g}\bar{f}g_{0})\,|g_{0}\rangle\langle g_{0}|$
hence
$\alpha_{\geq}^{(f)}\big{(}\operatorname{\Phi}_{0}(g,h)\big{)}\operatorname{\Phi}_{0}(f,gh)=\sum_{g_{0}\in G}\omega(g,h,\bar{h}\bar{g}\bar{f}g_{0})\omega(f,gh,\,\bar{h}\bar{g}\bar{f}g_{0})\,|g_{0}\rangle\langle g_{0}|$
Comparing this to
$\operatorname{\Phi}_{0}(f,g)\operatorname{\Phi}_{0}(fg,h)=\sum_{g_{0}\in G}\omega(f,g,\bar{g}\bar{f}g_{0})\omega(fg,h,\bar{h}\bar{g}\bar{f}g_{0})|g_{0}\rangle\langle g_{0}|$
and using the cocycle relation yields
$\operatorname{\Phi}_{0}(f,g)\operatorname{\Phi}_{0}(fg,h)=\omega(f,g,h)\times\alpha_{\geq}^{(f)}\big{(}\operatorname{\Phi}_{0}(g,h)\big{)}\operatorname{\Phi}_{0}(f,gh),$
showing that $\alpha$ indeed has anomaly $[\omega]$.
Appendix E Right restrictions and covariance
E.1 Right restrictions that are group homomorphisms
The following Lemma says that any symmetry with trivial anomaly admits right restrictions that are group homomorphisms, at least after stacking with an auxiliary gauge degree of freedom. The proof is closely analogous to [22, Theorem 3.1.6].
Lemma E.1.
Let $\alpha$ be a symmetry of range $R$ with trivial anomaly. For any site $j\in{\mathbb{Z}}$ we can enlarge the on-site algebra ${\mathcal{A}}_{j}$ by tensoring with a gauge degree of freedom $\operatorname{\mathrm{End}}({\mathbb{C}}^{\left|G\right|})$, obtaining an enlarged spin chain $\widetilde{\mathcal{A}}$. Then the symmetry $\tilde{\alpha}=\alpha\otimes\mathrm{id}$ on $\widetilde{\mathcal{A}}$ admits a right restriction $\tilde{\beta}_{\geq j}$ at site $j$ of defect size $5R$ such that $g\mapsto\tilde{\beta}_{\geq j}^{(g)}$ is a group homomorphism.
Proof.
Let $\alpha_{\geq j}$ be a right restriction of $\alpha$ at $j\in{\mathbb{Z}}$ with defect size $2R$. We drop $j$ from the notation in the remainder of this proof. Let $\operatorname{\Phi}(g,h)\in{\mathcal{A}}_{[j-2R,j+3R]}$ be the fusion operators associated to this right restriction. Since $\alpha$ has trivial anomaly, by Eq. C.1 we can choose the phases of the fusion operators such that
$\operatorname{\Phi}(f,g)\operatorname{\Phi}(fg,h)=\alpha_{\geq}^{(f)}\big{(}\operatorname{\Phi}(g,h)\big{)}\operatorname{\Phi}(f,gh)$ (E.1)
for all $f,g,h\in G$. Let $\tilde{\alpha}_{\geq}=\alpha_{\geq}\otimes\mathrm{id}$, which is a left restriction for $\tilde{\alpha}=\alpha\otimes\mathrm{id}$ of defect size $2R$. Define unitaries
$V(g):=\sum_{k}\operatorname{\Phi}(g,k)\otimes|k\rangle\langle gk|\in\widetilde{\mathcal{A}}_{[j-2R,j+3R]},$
then using Eq. (E.1) we obtain
$\tilde{\alpha}_{\geq}^{(g)}(V(h))V(g)V(gh)^{*}=\operatorname{\Phi}(g,h).$
Define a new right restriction $\tilde{\beta}_{\geq}$ of $\tilde{\alpha}$ with components $\tilde{\beta}_{\geq}^{(g)}=\mathrm{Ad}[V(g)^{*}]\circ\tilde{\alpha}_{\geq}^{(g)}$. Then
$\tilde{\beta}_{\geq}^{(g)}\circ\tilde{\beta}_{\geq}^{(h)}$ $=\mathrm{Ad}[V(g)^{*}]\circ\tilde{\alpha}_{\geq}^{(g)}\circ\mathrm{Ad}[V(h)^{*}]\circ\tilde{\alpha}_{\geq}^{(h)}$
$=\mathrm{Ad}[V(g)^{*}\,\tilde{\alpha}_{\geq}^{(g)}(V(h)^{*})\,\operatorname{\Phi}(g,h)]\circ\tilde{\alpha}_{\geq}^{(gh)}$
$=\mathrm{Ad}[V(gh)^{*}]\circ\tilde{\alpha}_{\geq}^{(gh)}=\tilde{\beta}_{\geq}^{(gh)}.$
i.e. $g\mapsto\tilde{\beta}_{\geq}^{(g)}$ is a group homomorphism. Finally noting that $\tilde{\beta}_{\geq}$ is a right restriction of $\tilde{\alpha}$ at $j$ with defect size $5R$ yields the claim. ∎
E.2 Covariant right restrictions
Let $\alpha$ be a symmetry of range $R$ on a spin chain ${\mathcal{A}}$, and let $\alpha_{\geq j}$ be a right restriction of $\alpha$ at some site $j$. We say $\alpha_{\geq j}$ is covariant if
$\alpha^{(k)}\circ\alpha_{\geq j}^{(\bar{k}gk)}\circ\alpha^{(\bar{k})}=\alpha_{\geq j}^{(g)}$
for all $g,k\in G$. The failure of the left restriction to be covariant is captured by local unitaries $\operatorname{\Psi}_{g}(k)$ which are uniquely defined up to phase by
$\alpha^{(k)}\circ\alpha_{\geq j}^{(\bar{k}gk)}\circ\alpha^{(\bar{k})}\circ\big{(}\alpha_{\geq j}^{(g)}\big{)}^{-1}=\mathrm{Ad}[\operatorname{\Psi}_{g}(k)].$
We call these the crossing operators associated to the right restriction $\alpha_{\geq j}$. If $\alpha_{\geq j}$ has defect size $L$ then $\operatorname{\Psi}_{g}(k)$ is supported on the interval $[j-(L+2R),j+(L+2R)]$.
By straightforward computation we find
$\mathrm{Ad}\left[\alpha^{(k)}\big{(}\operatorname{\Psi}_{\bar{k}gk}(l)\big{)}\,\operatorname{\Psi}_{g}(k)\right]=\mathrm{Ad}\left[\operatorname{\Psi}_{g}(kl)\right]$
so there are phases $\lambda_{g}(k,l)\in U(1)$ such that
$\alpha^{(k)}\big{(}\operatorname{\Psi}_{\bar{k}gk}(l)\big{)}\,\operatorname{\Psi}_{g}(k)=\lambda_{g}(k,l)\times\operatorname{\Psi}_{g}(kl)$
for all $g,k,l\in G$.
Proposition E.2.
The phases $\lambda_{g}(k,l)$ satisfy the twisted 2-cocycle equations
$1=\frac{\lambda_{g}(k,lm)\lambda_{\bar{k}gk}(l,m)}{\lambda_{g}(k,l)\lambda_{g}(kl,m)}$
for all $g,k,l,m\in G$. They therefore define a cohomology class $[\lambda]\in H^{2}(G,U(1)[G])$. (See Appendix G for the relevant definitions.)
The class $[\lambda]$ depends only on the symmetry $\alpha$, so we obtain a well defined map
$\Lambda:\operatorname{\mathsf{Sym}}_{G}\rightarrow H^{2}(G,U(1)[G])$
which we call the obstruction to covariant right restrictions.
For symmetries $\alpha$ and $\beta$ of range $R$ this obstruction satisfies
  1. 1.
    If $\alpha$ is decoupled then $\Lambda(\alpha)=[1]$, the identity element of $H^{2}(G,U(1)[G])$.
  2. 2.
    $\Lambda$ is locally computable: If $\alpha$ and $\beta$ act on the same spin chain ${\mathcal{A}}$ and there is an interval $I$ of length $112R+$ 1 such that $\alpha|_{{\mathcal{A}}_{I}}=\beta|_{{\mathcal{A}}_{I}}$ then $\Lambda(\alpha)=\Lambda(\beta)$.
  3. 3.
    $\Lambda$ is multiplicative under stacking: $\Lambda(\alpha\otimes\beta)=\Lambda(\alpha)\cdot\Lambda(\beta)$.
  4. 4.
    $\Lambda$ is constant on stable equivalence classes: $\alpha\sim\beta\,\implies\,\Lambda(\alpha)=\Lambda(\beta)$.
In particular, the obstruction lifts to a homomorphism of monoids $\Lambda:(\operatorname{\mathsf{Sym}}_{G}/\sim)\rightarrow H^{2}(G,U(1)[G])$.
Proof.
Can be found in Appendix I. ∎
Remark E.3.
  1. 1.
    It will follow from Theorem B.1 that $(\operatorname{\mathsf{Sym}}_{G}/\sim)$ is a group and so $\Lambda:(\operatorname{\mathsf{Sym}}_{G}/\sim)\rightarrow H^{2}(G,U(1)[G])$ is in fact a group homomorphism.
  2. 2.
    We could use left restrictions instead of right restrictions to define an obstruction to covariant left restrictions $\Lambda_{L}(\alpha)$. Then $\Lambda_{L}(\alpha)=\Lambda(\alpha)^{-1}$.
  3. 3.
    We will show in Appendix J that the obstruction $\Lambda(\alpha)$ is given by the inverse of the slant product of the anomaly of $\alpha$ (see Appendix G for definitions). In the context of SPTs the obstruction $\Lambda$ is therefore intimately related to the types and properties of the anyons supported by the gauged SPT [18], which are believed to be described by the Dijkgraaf-Witten TQFT [19] corresponding to the anomaly.
As the name suggests, if a symmetry $\alpha$ has vanishing obstruction to covariant right restrictions, then (after adding an ancilla) $\alpha$ indeed admits covariant right restrictions:
Lemma E.4.
Let $\alpha$ be a symmetry of range $R$ such that $\Lambda(\alpha)=[1]$. For any site $j\in{\mathbb{Z}}$ we can enlarge the on-site algebra ${\mathcal{A}}_{j}$ by tensoring with a gauge degree of freedom $\operatorname{\mathrm{End}}({\mathbb{C}}^{\left|G\right|})$, obtaining an enlarged spin chain $\widetilde{\mathcal{A}}$. Then the symmetry $\tilde{\alpha}=\alpha\otimes\rho_{\operatorname{\mathrm{reg}}}$, where $\rho_{\operatorname{\mathrm{reg}}}$ is the regular representation of $G$ on the gauge degree of freedom, admits a covariant right restriction at $j$ with defect size $5R$.
Proof.
Let $\alpha_{\geq}$ be a right restriction of $\alpha$ at $j\in{\mathbb{Z}}$ with defect size $2R$. Let $\operatorname{\Psi}_{g}(k)\in{\mathcal{A}}_{[j-4R,j+4R]}$ be the crossing operators associated to this right restriction. Since $\Lambda(\alpha)=[1]$ we can choose the phases of the $\operatorname{\Psi}_{g}(k)$ so that
$\alpha^{(k)}\big{(}\operatorname{\Psi}_{\bar{k}gk}(l)\big{)}\,\operatorname{\Psi}_{g}(k)=\operatorname{\Psi}_{g}(kl)$ (E.2)
for all $g,k,l\in G$.
Let $\tilde{\alpha}_{\geq}=(\alpha_{\geq}\otimes\mathrm{id})$, which is a right restriction of $\tilde{\alpha}=\alpha\otimes\rho_{\operatorname{\mathrm{reg}}}$ of defect size $2R$. Define unitaries
$V_{g}=\sum_{k}\,\operatorname{\Psi}_{g}(k)\otimes|k\rangle\langle k|\in\widetilde{\mathcal{A}}_{[j-4R,j+4R]}.$
Then
$V_{g}^{*}\,\tilde{\alpha}^{(k)}\big{(}V_{\bar{k}gk}\big{)}$ $=\left(\sum_{l_{1}}\,\operatorname{\Psi}_{g}(l_{1})^{*}\otimes|l_{1}\rangle\langle l_{1}|\right)\times\left(\sum_{l_{2}}\,\tilde{\alpha}^{(k)}\big{(}\operatorname{\Psi}_{\bar{k}gk}(k_{2})\big{)}\otimes|kl_{2}\rangle\langle kl_{2}|\right)$
putting $l=l_{1}=hl_{2}$ and using Eq. (E.2) this becomes
$=\sum_{l}\,\operatorname{\Psi}_{g}(l)^{*}\,\tilde{\alpha}^{(k)}\big{(}\operatorname{\Psi}_{\bar{k}gk}(\bar{k}l)\big{)}\otimes|l\rangle\langle l|=\sum_{l}\,\operatorname{\Psi}_{g}(k)^{*}\otimes|l\rangle\langle l|=\operatorname{\Psi}_{g}(k)^{*}.$
Now consider the right restriction $\tilde{\beta}_{\geq}$ with components $\tilde{\beta}_{\geq}^{(g)}=\mathrm{Ad}[V_{g}]\circ\tilde{\alpha}_{\geq}^{(g)}$. Then
$\tilde{\alpha}^{(k)}\circ\tilde{\beta}_{\geq}^{(\bar{k}gk)}\circ\tilde{\alpha}^{(\bar{k})}\circ\big{(}\tilde{\beta}_{\geq}^{(g)}\big{)}^{-1}$ $=\tilde{\alpha}^{(k)}\circ\mathrm{Ad}[V_{\bar{k}gk}]\circ\tilde{\alpha}_{\geq}^{(\bar{k}gk)}\circ\tilde{\alpha}^{(\bar{k})}\circ\big{(}\tilde{\alpha}_{\geq}^{(g)}\big{)}^{-1}\circ\mathrm{Ad}[V_{g}^{*}]$
$=\mathrm{Ad}\left[\tilde{\alpha}^{(k)}\big{(}V_{\bar{k}gk}\big{)}\operatorname{\Psi}_{g}(k)\,V_{g}^{*}\right]=\mathrm{id},$
so $\tilde{\beta}_{\geq}$ is covariant. Noting that $\tilde{\beta}_{\geq}$ is a right restriction of $\tilde{\alpha}$ at $j$ with defect size $5R$ finishes the proof. ∎
We now show that symmetries with trivial anomaly have no obstruction to covariant right restrictions.
Lemma E.5.
If $\alpha$ has trivial anomaly then $\Lambda(\alpha)=[1]$.
Proof.
Suppose $\alpha$ has range $R$. We can enlarge the on-site algebra ${\mathcal{A}}_{j}$ by tensoring with $\operatorname{\mathrm{End}}({\mathbb{C}}^{\left|G\right|})$, obtaining an extended spin chain $\widetilde{\mathcal{A}}$. By Lemma E.1 there is a right restriction $\tilde{\alpha}_{\geq}$ of $\tilde{\alpha}=\alpha\otimes\mathrm{id}$ at $j$ with defect size $5R$ and such that $g\mapsto\tilde{\alpha}_{\geq}^{(g)}$ is a group homomorphism. We can therefore regard $\tilde{\alpha}_{\geq}$ as a symmetry.
Since $\tilde{\alpha}_{\geq}$ and $\alpha$ agree everywhere to the right of the site $5R$ it follows from local computability of the obstruction to covariant right restrictions (item 2 of Proposition E.2) that $\Lambda(\alpha)=\Lambda(\tilde{\alpha}_{\geq})$. But $\tilde{\alpha}_{\geq}$ agrees with $\mathrm{id}$ everywhere to the left of the site $-5R$, so $\Lambda(\tilde{\alpha}_{\geq})=\Lambda(\mathrm{id})=[1]$ by local computability. We conclude that $\Lambda(\alpha)=[1]$. ∎
E.3 Covariant right restrictions that are group homomorphisms
Lemma E.6.
Let $\alpha$ be a symmetry of range $R$ with trivial anomaly on a spin chain ${\mathcal{A}}$. For any site $j\in{\mathbb{Z}}$ we can enlarge the on-site algebra ${\mathcal{A}}_{j}$ by tensoring with two gauge degrees of freedom $\operatorname{\mathrm{End}}({\mathbb{C}}^{\left|G\right|})\otimes\operatorname{\mathrm{End}}({\mathbb{C}}^{\left|G\right|})$, obtaining an enlarged spin chain ${\mathcal{A}}^{\prime\prime}$. There is an extension $\alpha=\alpha^{\prime\prime}\otimes\rho$ of $\alpha$ to the enlarged spin chain and a covariant right restriction $\beta_{\geq}$ of $\alpha^{\prime\prime}$ at $j$ with defect size $7R$ such that $g\mapsto\beta_{\geq}^{(g)}$ is a group homomorphism.
Proof.
For any site $j\in{\mathbb{Z}}$ we can enlarge the on-site algebra ${\mathcal{A}}_{j}$ by tensoring with a gauge degree of freedom $\operatorname{\mathrm{End}}({\mathbb{C}}^{\left|G\right|})$, obtaining an enlarged spin chain ${\mathcal{A}}^{\prime}$. Let $\alpha^{\prime}=\alpha\otimes\rho_{\operatorname{\mathrm{reg}}}$. Then Lemmas E.4 and E.5 imply that there is a covariant right restriction $\alpha^{\prime}_{\geq}$ of $\alpha^{\prime}$ at $j$ of defect size $5R$.
Since $\alpha^{\prime}$ is a symmetry of range $R$ with trivial anomaly we can repeat the construction in the proof of Lemma E.1 to obtain from $\alpha^{\prime}_{\geq}$ a new right restriction that is a group homomorphism. We point out here that this construction does not spoil the covariance.
To this end, let $\operatorname{\Phi}(g,h)\in{\mathcal{A}}_{[j-5R,j+7R]}$ be fusion operators associated with $\alpha^{\prime}_{\geq}$, with phases chosen such that
$\operatorname{\Phi}(f,g)\operatorname{\Phi}(fg,h)=(\alpha^{\prime}_{\geq})^{(f)}\big{(}\operatorname{\Phi}(g,h)\big{)}\operatorname{\Phi}(f,gh).$ (E.3)
Further enlarging the on-site algebra ${\mathcal{A}}^{\prime}_{j}$ by tensoring with another gauge degree of freedom $\operatorname{\mathrm{End}}({\mathbb{C}}^{\left|G\right|})$ we obtain an enlarged spin chain ${\mathcal{A}}^{\prime\prime}$. As in the proof of Lemma E.1 we define unitaries $V(g)=\sum_{k}\operatorname{\Phi}(g,k)\otimes|k\rangle\langle gk|$ supported on $[j-5R,j+6R]$ so that $\beta_{\geq}^{(g)}=\mathrm{Ad}[V(g)^{*}]\circ(\alpha^{\prime}_{\geq}\otimes\mathrm{id})$ defines a left restriction of $\alpha^{\prime}\otimes\mathrm{id}$ such that $g\mapsto\beta_{\geq}^{(g)}$ is a group homomorphism.
Covariance of $\alpha^{\prime}_{\geq}$ applied to the definition of the fusion operators results in
$(\alpha^{\prime})^{(k)}\big{(}\operatorname{\Phi}(g,h)\big{)}=\mu_{k}(g,h)\times\operatorname{\Phi}(kg\bar{k},kh\bar{k})$
for some phases $\mu_{k}(g,h)$ that satisfy
$\mu_{k}(lg\bar{l},lh\bar{l})\mu_{l}(g,h)=\mu_{kl}(g,h)\quad\text{and}\quad\mu_{k}(f,g)\mu_{k}(fg,h)=\mu_{k}(g,h)\mu_{k}(f,gh).$
The first of these relations follows from computing $(\alpha^{\prime})^{(kl)}\big{(}\operatorname{\Phi}(g,h)\big{)}$ in two different ways. The second relation, which says that the $\mu_{k}$ are 2-cocycles, follows from (E.3).
Using these relations one easily verifies that
$\rho_{\mu}^{(k)}(|f\rangle\langle g|)=\frac{\mu_{k}(g,\bar{f})}{\mu_{k}(\bar{f},f)\mu_{k}(g,f)}|kf\bar{k}\rangle\langle kg\bar{k}|$
defines a unitary $G$-action on the gauge degree of freedom. Taking $\alpha^{\prime\prime}=\alpha^{\prime}\otimes\rho_{\mu}$ we find
$(\alpha^{\prime\prime})^{(k)}\big{(}V(g)\big{)}=\sum_{f\in G}\,\frac{\mu_{k}(g,f)\mu_{k}(gf,\bar{f})}{\mu_{k}(\bar{f},f)\mu_{k}(gf,f)}\operatorname{\Phi}(kg\bar{k},kf\bar{k})\otimes|kf\bar{k}\rangle\langle kg\bar{k}kf\bar{k}|=V(kg\bar{k})$
where we used the 2-cocycle relation for $\mu_{k}$ on elements $(gf,\bar{f},f)$.
This implies that $\beta_{\geq}^{(g)}=\mathrm{Ad}[V(g)^{*}]\circ(\alpha^{\prime}_{\geq}\otimes\mathrm{id})$ is a covariant right restriction of $\alpha^{\prime\prime}$ at $j$ with defect size $7R$. ∎
Appendix F Proof of the main Theorem
Proposition F.1.
Let $\alpha$ be a symmetry with trivial anomaly. Then $\alpha$ is stably equivalent to a decoupled symmetry.
Proof.
We will decouple the symmetry by splitting $\alpha$ into local components which commute with each other. We then construct an explicit equivalence with a block partitioned QCA.
Let $R$ be the range of $\alpha$ and take $L=16R$. By repeated use of Lemma E.6 we can stack with a decoupled symmetry $\alpha_{\operatorname{\mathrm{triv}}}=\bigotimes_{j\in L{\mathbb{Z}}}\,\rho_{j}$, where $\rho_{j}$ acts on site $j$ only, to obtain the symmetry $\alpha^{\prime}=\alpha\otimes\alpha_{\operatorname{\mathrm{triv}}}\sim\alpha$ of range $R$ on a stacked spin chain ${\mathcal{A}}^{\prime}$ such that $\alpha^{\prime}$ admits covariant right restrictions $\alpha_{\geq j}^{\prime}$ of defect size $7R$ at all sites $j\in L{\mathbb{Z}}$ and such that $g\mapsto\alpha_{\geq j}^{\prime(g)}$ are group homomorphisms.
Covariance implies that whenever $j\geq i+L$ we have
$\alpha_{\geq i}^{\prime(k)}\circ\alpha_{\geq j}^{\prime(\bar{k}gk)}\circ\alpha_{\geq i}^{\prime(\bar{k})}=\alpha^{\prime(k)}\circ\alpha_{\geq j}^{\prime(\bar{k}gk)}\circ\alpha^{\prime(\bar{k})}=\alpha_{\geq j}^{\prime(g)}.$ (F.1)
For each $j\in{\mathbb{Z}}$, define $\alpha_{j}^{\prime(g)}:=\alpha_{\geq jL}^{\prime(g)}\circ(\alpha_{\geq(j+1)L}^{\prime(g)})^{-1}$. Then it follows from Eq. (F.1) that
  1. 1.
    $[\alpha_{i}^{\prime(g)},\alpha_{j}^{\prime(h)}]=0$ whenever $i\neq j$.
  2. 2.
    $g\mapsto\alpha_{j}^{\prime(g)}$ are group homomorphisms.
  3. 3.
    $\alpha^{\prime}=\prod_{j\in{\mathbb{Z}}}\alpha^{\prime}_{j}$.
  4. 4.
    $\alpha^{\prime}_{j}$ is supported on $[jL-7R,(j+1)L+8R]$.
Pick unitaries $U_{j}(g)\in{\mathcal{A}}^{\prime}$ such that $\alpha_{j}^{\prime(g)}=\mathrm{Ad}[U_{j}(g)]$. Since $g\mapsto\alpha_{j}^{\prime(g)}$ is a group homomorphism, the unitaries $U_{j}(g)$ form a projective representation of $G$. The unitaries $U_{j}(g)$ commute with unitaries $U_{j^{\prime}}(h)$ whenever $\left|j^{\prime}-j\right|\geq 2$ because they have disjoint supports. Since $\alpha_{j-1}^{\prime(g)}$ commutes with $\alpha_{j}^{\prime(h)}$ we find that there are $\chi_{j}:G^{2}\rightarrow U(1)$ such that
$U_{j-1}(g)U_{j}(h)=\chi_{j}(g,h)\,U_{j}(h)U_{j-1}(g).$ (F.2)
Our first goal is now to upgrade these unitaries $U_{j}$ to unitaries $\widetilde{U}_{j}$ with the same spatial support, and such that $g\mapsto\widetilde{U}_{j}(g)$ are still projective representations, but also such that
$[\widetilde{U}_{j}(g),\widetilde{U}_{i}(h)]=0,\qquad i\neq j,$ (F.3)
instead of (F.2).
In order to achieve this, we need to stack again. At each site $jL$ with $j\in{\mathbb{Z}}$, we add a degree of freedom with algebra ${\mathcal{C}}_{j}$ which is a copy of ${\mathcal{A}}_{[L(j-1),L(j+1)]}$, obtaining the enlarged spin chain ${\mathcal{A}}^{\prime\prime}$. We recall that the projective representations $U_{j-1}$ and $U_{j}$ map in ${\mathcal{A}}_{[L(j-1),L(j+1)]}$. Let us now consider the conjugate representations $\overline{U}_{j-1}$ and $\overline{U}_{j}$ mapping in the copy ${\mathcal{C}}_{j}$. The conjugate representations will always be considered inside ${\mathcal{C}}_{j}$ whereas the original representations act on the spin chain ${\mathcal{A}}^{\prime}$. We now define the symmetry $\rho_{Lj}^{\prime\prime}$ on the added degrees of freedom in ${\mathcal{C}}_{j}$ given by
$(\rho^{\prime\prime}_{jL})^{(g)}=\mathrm{Ad}[\overline{U}_{j-1}(g)\overline{U}_{j}(g)]$
To check that this is indeed a representation, we recall that $U_{j-1},U_{j}$, and hence also $\overline{U}_{j-1},\overline{U}_{j}$, commute up to a phase. In fact, we have
$\overline{U}_{j-1}(g)\overline{U}_{j}(h)=\bar{\chi}_{j}(g,h)\,\overline{U}_{j}(h)\overline{U}_{j-1}(g).$ (F.4)
Writing $\rho^{\prime\prime}=\bigotimes_{j\in{\mathbb{Z}}}\,\rho^{\prime\prime}_{jL}$ we thus obtain a new symmetry $\alpha^{\prime\prime}=\alpha^{\prime}\otimes\rho^{\prime\prime}\sim\alpha^{\prime}$ acting on the new spin chain ${\mathcal{A}}^{\prime\prime}$. Finally, we define
$\widetilde{U}_{j}(g)=\overline{U}_{j}(g)\otimes U_{j}(g)\otimes\overline{U}_{j}(g)\in{\mathcal{C}}_{j}\otimes{\mathcal{A}}^{\prime}\otimes{\mathcal{C}}_{j+1}$
We note that $\widetilde{U}_{j}$ is a projective representation since it is a tensor product of projective representations. Writing $\alpha_{j}^{\prime\prime(g)}:=\mathrm{Ad}[\widetilde{U}_{j}(g)]$, we can verify that all properties of $\alpha_{j}^{\prime}$ listed below (F.1) hold for $\alpha_{j}^{\prime\prime}$ as well, in particular, $\alpha^{\prime\prime}=\bigotimes_{j\in{\mathbb{Z}}}\alpha_{j}^{\prime\prime}$. However, as announced, we now have the stronger commutation property (F.3). This is checked by an explicit computation using (F.2), (F.4) and $\chi_{j}(g,h)\bar{\chi}_{j}(g,h)=1$.
We now proceed to decouple $\alpha^{\prime\prime}$. This requires more stacking. For each $j\in{\mathbb{Z}}$ let $m(j)=jL+8R$ be the site sitting half way between $jL$ and $(j+1)L$. Stack with a spin chain ${\mathcal{B}}$ with on-site algebras ${\mathcal{B}}_{m(j)}=\operatorname{\mathrm{End}}({\mathbb{C}}^{\left|G\right|})$ for all $j$ and ${\mathcal{B}}_{j}\simeq{\mathbb{C}}$ at all other sites to obtain a new spin chain ${\mathcal{A}}^{\prime\prime\prime}$. Let $\rho^{\prime\prime\prime}_{\operatorname{\mathrm{reg}}}$ be the decoupled symmetry on ${\mathcal{B}}$ which acts with the regular representation $\rho^{\prime\prime\prime}_{\operatorname{\mathrm{reg}}}$ on each ${\mathcal{B}}_{m(j)}$. Let $\alpha^{\prime\prime\prime}=\alpha^{\prime\prime}\otimes\rho^{\prime\prime\prime}_{\operatorname{\mathrm{reg}}}\sim\alpha^{\prime\prime}\sim\alpha$. We now explicitly construct a FDQC $\gamma$ such that $\gamma^{-1}\circ\alpha^{\prime\prime\prime}\circ\gamma$ is block partitioned.
Define unitaries
$V_{j}:=\sum_{h\in G}\,\widetilde{U}_{j}(h)\otimes|h\rangle\langle h|$
where the second tensor factor corresponds to ${\mathcal{B}}_{m(j)}$. Let $\gamma_{j}=\mathrm{Ad}[V_{j}]$, then $\gamma_{j}$ has the same support as $\alpha^{\prime\prime}_{j}$. Since $\widetilde{U}_{j}$ is a projective representation we have
$W_{j}(g):=V_{j}^{*}\,\big{(}\widetilde{U}_{j}(g)\otimes L_{m(j)}^{(g)}\big{)}\,V_{j}=\sum_{h\in G}\widetilde{U}_{j}(gh)^{*}\widetilde{U}_{j}(g)\widetilde{U}_{j}(h)\otimes|gh\rangle\langle g|=\sum_{h\in G}\,\operatorname{\mathbbm{1}}\,\otimes\,c_{j}(g,h)|gh\rangle\langle h|,$
for some 2-cocycle $c_{j}:G^{2}\rightarrow U(1)$. Here $L_{m(j)}^{(g)}\in{\mathcal{B}}_{m(j)}$ is the left regular action. The unitary $W_{j}(g)$ is seen to act non-trivially only on ${\mathcal{B}}_{m(j)}$ and so
$\gamma_{j}^{-1}\circ\big{(}\alpha_{j}^{\prime\prime(g)}\,\otimes\,\rho_{\operatorname{\mathrm{reg}}}^{\prime\prime\prime(g)}\big{)}\otimes\gamma_{j}=\mathrm{Ad}[W_{j}(g)]$
is a $G$-action which acts non-trivially only on ${\mathcal{B}}_{m(j)}$. Moreover, from Eq. (F.3) we find that each $\gamma_{j}$ commutes with every component of $\alpha_{i}^{\prime\prime}$ whenever $i\neq j$. Defining the FDQC $\gamma=\prod_{j\in{\mathbb{Z}}}\gamma_{j}$ and recalling that $\alpha^{\prime\prime}=\prod_{j\in{\mathbb{Z}}}\alpha^{\prime\prime}_{j}$ we therefore find that
$\beta^{(g)}:=\gamma^{-1}\circ\alpha^{\prime\prime\prime(g)}\circ\gamma=\gamma^{-1}\circ(\alpha^{\prime\prime(g)}\otimes\rho_{\operatorname{\mathrm{reg}}}^{\prime\prime\prime(g)})\circ\gamma=\prod_{j\in{\mathbb{Z}}}\mathrm{Ad}[W_{j}(g)].$
We conclude that $\alpha\sim\beta=\bigotimes_{j\in{\mathbb{Z}}}\,\mathrm{Ad}[W_{j}]$ is a decoupled symmetry. ∎
Corollary F.2.
Let $\alpha$ and $\beta$ be symmetries on spin chains ${\mathcal{A}}$ and ${\mathcal{B}}$ respectively. If $\alpha$ and $\beta$ have the same anomaly then they are stably equivalent.
Proof.
Let $\gamma$ be a symmetry with anomaly $\operatorname{\Omega}(\gamma)=\operatorname{\Omega}(\alpha)^{-1}$, defined on a spin chain ${\mathcal{C}}$. (One could take the symmetry constructed in Section D).
Using multiplicativity of the anomaly under stacking and Proposition F.1 we find
$\alpha\sim\alpha\otimes(\gamma\otimes\beta)=(\alpha\otimes\gamma)\otimes\beta\sim\beta$
as required. ∎
We are now ready to prove our main Theorem B.1.
Proof of Theorem B.1 : To show that the monoid $(\operatorname{\mathsf{Sym}}_{G}/\sim)$ is in fact a group it suffices to show that each class has an inverse. Let $\alpha$ be a symmetry. Then by Section D there exists a symmetry $\beta$ such that $\Omega(\beta)=\Omega(\alpha)^{-1}$. Since the anomaly is multiplicative under stacking we have that $\alpha\otimes\beta$ has trivial anomaly. By Proposition F.1 the stacked symmetry $\alpha\otimes\beta$ is stably equivalent to a decoupled symmetry. This shows that the stable equivalence class of $\beta$ is the inverse of the class of $\alpha$ in $(\operatorname{\mathsf{Sym}}_{G}/\sim)$.
The statement that two symmetries are stably equivalent if and only if their anomalies are equal follows from item 2 of Proposition C.2 and Corollary F.2. The fact that $\Omega$ lifts to a group homomorphism from $(\operatorname{\mathsf{Sym}}_{G}/\sim)$ to $H^{3}(G,U(1))$ also follows from Proposition C.2. That this homomorphism is actually an isomorphism follows from the examples of Section D, which provide for each $[\omega]\in H^{3}(G,U(1))$ a symmetry $\alpha$ with $\operatorname{\Omega}(\alpha)=[\omega]$.
[…]
Appendix G (Twisted) group cohomology
We give the necessary definitions of (twisted) group cohomology and the slant product. For an in depth treatment, see for example the monograph [23].
Let $G$ be a group and $M$ a $G$-module. Consider the co-chain complex
[…]
where $C^{n}$ is the collection of all functions $G^{n}\to M$ and $\delta$ is the differential map which sends a function $C^{n}\ni\theta:G^{n}\to M$ to $\delta\theta\in C^{n+1}$ defined by
$(\delta\theta)(g_{1},...,g_{n+1})=g_{1}\cdot\theta(g_{2},...,g_{n+1})\left(\prod\limits_{j=1}^{n}\theta(g_{1},...,g_{i}g_{i+1},...,g_{n+1})^{(-1)^{j}}\right)\theta(g_{1},...,g_{n})^{(-1)^{n+1}}.$
The map $\delta$ is called the co-boundary map, it satisfies $\delta^{2}=\mathrm{id}$.
The $n$-cocycles $Z^{n}$ are the functions in $C^{n}$ in the kernel of the co-boundary map $\delta$. The $n$-coboundaries $B^{n}$ are the functions of $C^{n}$ in the image of $\delta:C^{n-1}\to C^{n}$. The $n$-th cohomology group $H^{n}(G,M)$ is defined by $H^{n}(G,M)=Z^{n}/B^{n}$ together with the multiplication $[\theta]\cdot[\phi]=[\theta\phi]$ where $(\theta\phi)(g_{1},...,g_{n})=\theta(g_{1},...,g_{n})\cdot\phi(g_{1},...,g_{n})$ for two representatives $\theta,\phi\in Z^{n}$. The quotient is well defined because $\delta^{2}=1$.
In the following, and throughout the rest of this paper, the words cocycle, coboundary and cohomology will always refer to the case $M=U(1)$ with the trivial action of $G$ on $U(1)$. The words twisted cocycle, twisted coboundary and twisted cohomology will always refer to the case $M=U(1)[G]$, the collection of functions $G\to U(1)$ with the following action. For $C^{n}\ni\theta$ write $\theta_{h}$ for the assignment $\theta:G^{n}\to U(1)[G]:(g_{1},...,g_{n})\mapsto(h\mapsto\theta_{h}(g_{1},...,g_{n}))$. The action of $f\in G$ on $\theta$ is defined by
$(f\cdot\theta)(g_{1},...,g_{n})=(h\mapsto\theta_{\bar{f}hf}(g_{1},...,g_{n}))$
or pointwise by $f\cdot\theta_{h}=\theta_{\bar{f}hf}$.
A map $\omega:G^{3}\to U(1)$ is a (normal) 3-cocycle if
$\delta\omega(g_{1},g_{2},g_{3},g_{4})=\frac{\omega(g_{2},g_{3},g_{4})\omega(g_{1},g_{2}g_{3},g_{4})\omega(g_{1},g_{2},g_{3})}{\omega(g_{1}g_{2},g_{3},g_{4}),\omega(g_{1},g_{2},g_{3}g_{4})}=1.$ (G.1)
Two 3-cocycles are equivalent (represent the same cohomology class) if they are equal up to multiplication by a 3-coboundary
$(g_{1},g_{2},g_{3})\mapsto\frac{\xi(g_{2},g_{3})\xi(g_{1},g_{2}g_{3})}{\xi(g_{1}g_{2},g_{3})\xi(g_{1},g_{2})}$ (G.2)
for some $\xi:G^{2}\to U(1)$.
A map $\lambda:G^{2}\to U(1)[G]$ is a twisted 2-cocycle if
$(\delta\lambda)(g_{1},g_{2},g_{3})=h\mapsto\frac{\lambda_{\bar{g}_{1}hg_{1}}(g_{2},g_{3})\lambda_{h}(g_{1},g_{2}g_{3})}{\lambda_{h}(g_{1}g_{2},g_{3})\lambda_{h}(g_{1},g_{2})}\equiv 1.$ (G.3)
Two twisted 2-cocycles are equivalent (represent the same twisted cohomology class) if they are equal up to multiplication by a twisted 2-coboundary
$h\mapsto\frac{\epsilon_{\bar{g}_{1}hg_{1}}(g_{2})\epsilon_{h}(g_{1})}{\epsilon_{h}(g_{1}g_{2})}$ (G.4)
for some $\epsilon:G\to U(1)[G]$.
Given a 3-cocycle $\omega$ one obtains a twisted 2-cocycle $\tau(\omega)$ defined by
$\tau(\omega)_{g}(k,l)=\frac{\omega(g,k,l)\,\omega(k,l,\bar{l}\,\bar{k}gkl)}{\omega(k,\bar{k}gk,l)}$ (G.5)
for all $g,k,l\in G$. This lifts to a well defined group homomorphism $\tau:H^{3}(G,U(1))\rightarrow H^{2}(G,U(1)[G])$ called the slant product (also called the loop transgression).
Appendix H Proof of Proposition C.2
That the phases $\omega_{j}$ form a 3-cocycle, as well as the fact that the class $[\omega_{j}]$ is independent of the choice of right restriction and the choice of fusion operators is well known, see for example [17, Appendix B] for proofs. Since a right restriction at $j^{\prime}$ can be viewed as a right restriction at $j$ with perhaps a different defect size, this also shows independence from $j$. It follows that $\operatorname{\Omega}(\alpha)$ is well defined.
Let us now prove items 1 through 4.
  1. 1.
    If $\alpha$ is decoupled then we can take a right restriction $\alpha_{\geq}$ such that $g\mapsto\alpha_{\geq}^{(g)}$ is a group homomorphism. The associated fusion operators can all be taken to be the identity so the associated 3-cocycle is identically one. The anomaly is therefore trivial.
  2. 2.
    If $I=[a,b]$ is an interval of length $18R+$ 1 such that $\alpha|_{{\mathcal{A}}_{I}}=\beta|_{{\mathcal{A}}_{I}}$ then there is a site $j\in[a+4R,b-4R]$ and right restrictions $\alpha_{\geq}$ and $\beta_{\geq}$ of defect size $2R$ at $j$ such that $\alpha_{\geq}|_{{\mathcal{A}}_{I}}=\beta_{\geq}|_{{\mathcal{A}}_{I}}$. Let $\operatorname{\Phi}_{\alpha}(g,h)$ and $\operatorname{\Phi}_{\beta}(g,h)$ be fusion operators associated to these right restrictions. Since the automorphisms
    $\mathrm{Ad}[\operatorname{\Phi}_{\alpha}(g,h)]$ $=\alpha_{\geq}^{(g)}\circ\alpha_{\geq j}^{(h)}\circ\big{(}\alpha_{\geq j}^{(gh)}\big{)}^{-1}$
    $\mathrm{Ad}[\operatorname{\Phi}_{\beta}(g,h)]$ $=\beta_{\geq}^{(g)}\circ\beta_{\geq j}^{(h)}\circ\big{(}\beta_{\geq j}^{(gh)}\big{)}^{-1}$
    agree on ${\mathcal{A}}_{[a+2R,b-2R]}$ and the fusion operators are supported on $[a+2R,b-2R]$, it follows that $\operatorname{\Phi}_{\alpha}(g,h)$ and $\operatorname{\Phi}_{\beta}(g,h)$ are equal up to phases for all $g,h\in G$. It follows that the associated 3-cocycles are equal up to a coboundary and therefore represent the same element of $H^{3}(G,U(1))$.
  3. 3.
    Let $\alpha_{\geq}$ and $\beta_{\geq}$ be right restrictions at some $j\in{\mathbb{Z}}$ of $\alpha$ and $\beta$ respectively. Let $\operatorname{\Phi}_{\alpha}(g,h)$ and $\operatorname{\Phi}_{\beta}(g,h)$ be associated fusion operators and $\omega_{\alpha}$ and $\omega_{\beta}$ the corresponding 3-cocycles. Then $\alpha_{\geq}\otimes\beta_{\geq}$ is a right restriction of $\alpha\otimes\beta$ with associated fusion operators $\operatorname{\Phi}(g,h)=\operatorname{\Phi}_{\alpha}(g,h)\otimes\operatorname{\Phi}_{\beta}(g,h)$ and corresponding 3-cocycle $\omega=\omega_{\alpha}\cdot\omega_{\beta}$. Therefore $\operatorname{\Omega}(\alpha\otimes\beta)=[\omega]=[\omega_{\alpha}\cdot\omega_{\beta}]=[\omega_{\alpha}]\cdot[\omega_{\beta}]=\operatorname{\Omega}(\alpha)\cdot\operatorname{\Omega}(\beta)$, as required.
  4. 4.
    Suppose $\alpha\sim_{0}\beta$ are symmetries whose ranges are bounded by $R$, defined on the same spin chain ${\mathcal{A}}$. Then there is a FDQC $\gamma$ such that $\beta=\gamma^{-1}\circ\alpha\circ\gamma$. Since $\gamma$ is a FDQC there is a $C>0$ and a decomposition $\gamma=\gamma_{L}\circ\gamma_{R}$ of $\gamma$ in to FDQCs $\gamma_{L}$ and $\gamma_{R}$ such that $\gamma_{L}$ acts as identity on ${\mathcal{A}}_{\geq C}$ and $\gamma_{R}$ acts as identity on ${\mathcal{A}}_{\leq-C}$. Then
    $\alpha\sim_{0}\gamma_{L}^{-1}\circ\alpha\circ\gamma_{L}\sim_{0}\gamma_{R}^{-1}\circ\gamma_{L}^{-1}\circ\alpha\circ\gamma_{L}\circ\gamma_{R}=\beta.$
    But $\alpha$ and $\gamma_{L}^{-1}\circ\alpha\circ\gamma_{L}$ agree on ${\mathcal{A}}_{\geq(C+R)}$ so by local computability $\operatorname{\Omega}(\alpha)=\operatorname{\Omega}(\gamma_{L}^{-1}\circ\alpha\circ\gamma_{L})$. Similarly $\gamma_{L}^{-1}\circ\alpha\circ\gamma_{L}$ and $\beta$ agree on ${\mathcal{A}}_{\leq-(C+R)}$ so by local computability $\operatorname{\Omega}(\gamma_{L}^{-1}\circ\alpha\circ\gamma_{L})=\operatorname{\Omega}(\beta)$, yielding $\operatorname{\Omega}(\alpha)=\operatorname{\Omega}(\beta)$.
    If $\alpha^{\prime}$ is a decoupled symmetry then $\operatorname{\Omega}(\alpha^{\prime})=[1]$ by item 1 and $\operatorname{\Omega}(\alpha\otimes\alpha^{\prime})=\operatorname{\Omega}(\alpha)\cdot\operatorname{\Omega}{\alpha^{\prime}}=\operatorname{\Omega}(\alpha)$ by item 3. Together with the invariance of the anomaly under $\sim_{0}$, this shows that the anomaly is constant on stable equivalence classes.
Appendix I Proof of Proposition E.2
We first show that the class $[\lambda]\in H^{2}(G,U(1)[G])$ is independent of the choice of right restriction. Let $\alpha_{\geq}$ and $\tilde{\alpha}_{\geq}$ be right restriction of the symmetry $\alpha$. Then there are local unitaries $\{\Gamma_{g}\}_{g\in G}$ such that
$\tilde{\alpha}_{L}^{(g)}:=\mathrm{Ad}[\Gamma_{g}]\circ\alpha_{L}^{(g)}.$
If $\operatorname{\Psi}_{g}(k)$ are crossing operators associated to $\alpha_{\geq}$ then crossing operators $\tilde{\operatorname{\Psi}}_{g}(k)$ associated to $\tilde{\alpha}_{\geq}$ must satisfy
$\mathrm{Ad}[\tilde{\operatorname{\Psi}}_{g}(h)]=\alpha^{(h)}\circ\tilde{\alpha}_{L}^{(\bar{h}gh)}\circ\left(\alpha^{(h)}\right)^{-1}\circ\left(\tilde{\alpha}_{L}^{(g)}\right)^{-1}=\mathrm{Ad}\left[\alpha^{(h)}(\Gamma_{\bar{h}gh})\,\operatorname{\Psi}_{g}(h)\,\Gamma_{g}^{*}\right].$
Therefore
$\tilde{\operatorname{\Psi}}_{g}(h)=\varepsilon_{g}(h)\alpha^{(h)}\big{(}\Gamma_{\bar{h}gh}\big{)}\,\operatorname{\Psi}_{g}(h)\,\Gamma_{g}^{*}$ (I.1)
for some phase map $\varepsilon:G\to U(1)[G]$. Let $\lambda$ be the twisted 2-cocycle corresponding to the crossing operators $\operatorname{\Psi}_{g}(k)$. By a straightforward computation we find that the twisted 2-cocycle $\tilde{\lambda}$ corresponding to the $\tilde{\operatorname{\Psi}}_{g}(k)$ is
$\tilde{\lambda}_{g}(k,l)=\dfrac{\varepsilon_{\bar{k}gk}(l)\varepsilon_{g}(k)}{\varepsilon_{g}(kl)}\lambda_{g}(k,l).$
They differ up to a twisted 2-coboundary, conform (G.4), so $[\tilde{\lambda}]=[\lambda]$111 This construction immediately yields a way to turn $\lambda$’s which are coboundaries to 1, by setting $\tilde{\operatorname{\Psi}}(g,h)=\epsilon_{g}(h)\operatorname{\Psi}_{g}(h)$. . This shows that $\Lambda(\alpha)$ is well defined.
The proofs of items 1 through 4 are virtually identical to the corresponding proofs of items 1 through 4 in Appendix H. We do not repeat the details here.
Appendix J Computation of obstructions to covariant right restrictions
Let $\alpha$ be the symmetry with anomaly $[\omega]$ constructed in Section D. Let $\alpha_{\geq}$ be the right restriction at $a\in{\mathbb{Z}}$ consisting of all gates defining $\alpha$ that are supported in $[a,\infty)$. By decomposing
$\alpha^{(k)}=\alpha_{\leq}^{(k)}\circ\alpha_{\geq}^{(k)}\circ\mathrm{Ad}[U^{(k)}_{a-1,a}]$
we find that the associated crossing operators satisfy
$\mathrm{Ad}[\operatorname{\Psi}_{g}(k)]$ $=\alpha^{(k)}\circ\alpha_{\geq}^{(\bar{k}gk)}\circ\alpha^{(\bar{k})}\circ\big{(}\alpha_{\geq}^{(g)}\big{)}^{-1}$
$=\mathrm{Ad}[L_{a-1}^{(k)}]\circ\alpha_{\geq}^{(k)}\circ\mathrm{Ad}[U_{a-1,a}^{(k)}]\circ\alpha_{\geq}^{(\bar{k}gk)}\circ\mathrm{Ad}[U_{a-1,a}^{(k)}]^{*}\circ\big{(}\alpha_{\geq}^{(g)}\circ\alpha_{\geq}^{(k)}\big{)}^{-1}\circ\mathrm{Ad}[L_{a-1}^{(k)}]^{*}$
$=\mathrm{Ad}[L_{a-1}^{(k)}]\circ\alpha_{\geq}^{(k)}\circ\left(\mathrm{Ad}[U_{a-1,a}^{(k)}]\circ\alpha_{\geq}^{(\bar{k}gk)}\circ\mathrm{Ad}[U_{a-1,a}^{(k)}]^{*}\circ\big{(}\alpha_{\geq}^{(\bar{k}gk)}\big{)}^{-1}\right)$
$\quad\quad\quad\circ\alpha_{\geq}^{(\bar{k}gk)}\circ\big{(}\mathrm{Ad}[\operatorname{\Phi}_{a}(g,k)]\circ\alpha_{\geq}^{(gk)}\big{)}^{-1}\circ\mathrm{Ad}[L_{a-1}^{(k)}]^{*}$
where $\operatorname{\Phi}_{a}$ are the fusion operators associated to $\alpha_{\geq}$ as in Lemma D.1. The commutator expression is given by
$\mathrm{Ad}[U_{a-1,a}^{(k)}]\circ\alpha_{\geq}^{(\bar{k}gk)}\circ\mathrm{Ad}[U_{a-1,a}^{(k)}]^{*}\circ\big{(}\alpha_{\geq}^{(\bar{k}gk)}\big{)}^{-1}=\mathrm{Ad}[W_{a}(g,k)]$
where the unitary $W_{a}(g,k)$ is supported on $\{a-1,a\}$ and is diagonal in the group basis:
$W_{a}(g,k)|g_{a-1},g_{a}\rangle=\frac{\omega(k,g_{a},\bar{g}_{a}g_{a-1})}{\omega(k,\bar{k}\bar{g}kg_{a},\bar{g}_{a}\bar{k}gkg_{a-1})}\,|g_{a-1},g_{a}\rangle.$
Commuting $W_{a}(g,k)$ with $\mathrm{Ad}[L_{a-1}^{(k)}]\circ\alpha_{\geq}^{(k)}$ we obtain
$\mathrm{Ad}[\operatorname{\Psi}_{g}(k)]$ $=\mathrm{Ad}[W^{\prime}_{a}(g,k)]\circ\alpha_{\geq}^{(k)}\circ\alpha_{\geq}^{(\bar{k}gk)}\circ\big{(}\alpha_{\geq}^{(gk)}\big{)}^{-1}\circ\mathrm{Ad}[\operatorname{\Phi}_{a}(g,k)^{*}]$
$=\mathrm{Ad}[W^{\prime}_{a}(g,k)]\circ\mathrm{Ad}[\operatorname{\Phi}_{a}(k,\bar{k}gk)]\circ\mathrm{Ad}[\operatorname{\Phi}_{a}(g,k)^{*}]$
$=\mathrm{Ad}\big{[}W^{\prime}_{a}(g,k)\operatorname{\Phi}_{a}(k,\bar{k}gk)\operatorname{\Phi}(g,k)^{*}\big{]}$
with
$W^{\prime}_{a}(g,k)|g_{a-1},g_{a}\rangle=\frac{\omega(k,\bar{k}g_{a},\bar{g}_{a}g_{a-1})}{\omega(k,\bar{k}\bar{g}g_{a},\bar{g}_{a}gg_{a-1})}\,|g_{a-1},g_{a}\rangle.$
We can therefore take $\operatorname{\Psi}_{g}(k)=W_{a}(g,k)\operatorname{\Phi}_{a}(k,\bar{k}gk)\operatorname{\Phi}(g,k)^{*}$, which is supported on $\{a-1,a\}$.
Let us now compute the action of $\alpha^{(k)}\big{(}\operatorname{\Psi}_{\bar{k}gk}(l)\big{)}\,\operatorname{\Psi}_{g}(k)\operatorname{\Psi}_{g}(kl)^{*}$ on a product state $|(g_{i})\rangle$ in the group basis. Noting that $\alpha^{(k)}$ makes $\operatorname{\Psi}_{\bar{k}gk}(l)$ act on the product state $|(\bar{k}g_{i})\rangle$, we find
$\alpha^{(k)}$ $\big{(}\operatorname{\Psi}_{\bar{k}gk}(l)\big{)}\,\operatorname{\Psi}_{g}(k)\operatorname{\Psi}_{g}(kl)^{*}|(g_{i})\rangle=\left(\frac{\omega(l,\bar{l}\,\bar{k}g_{a},\bar{g}_{a}g_{a-1})\,\omega(l,\bar{l}\,\bar{k}gkl,\bar{l}\,\bar{k}\,\bar{g}g_{a})}{\omega(l,\bar{l}\,\bar{k}\,\bar{g}g_{a},\bar{g}_{a}gg_{a-1})\,\omega(\bar{k}gk,l,\bar{l}\,\bar{k}\,\bar{g}g_{a})}\right)$
×$\times\left(\frac{\omega(k,\bar{k}g_{a},\bar{g}_{a}g_{a-1})\,\omega(k,\bar{k}gk,\bar{k}\,\bar{g}g_{a})}{\omega(k,\bar{k}\,\bar{g}g_{a},\bar{g}_{a}gg_{a-1})\,\omega(g,k,\bar{k}\,\bar{g}\,g_{a})}\right)\times\left(\frac{\omega(kl,\bar{l}\,\bar{k}g_{a},\bar{g}_{a}g_{a-1})\,\omega(kl,\bar{l}\,\bar{k}gkl,\bar{l}\,\bar{k}\,\bar{g}g_{a})}{\omega(kl,\bar{l}\,\bar{k}\,\bar{g}g_{a},\bar{g}_{a}gg_{a-1})\,\omega(g,kl,\bar{l}\,\bar{k}\,\bar{g}g_{a})}\right)^{-1}\,|(g_{i})\rangle$
Successively applying 3-cocycle relations for elements $(k,l,\bar{l}\,\bar{k}g_{a},\bar{g}_{a}g_{a-1})$, then $(k,l,\bar{l}\,\bar{k}\,\bar{g}g_{a},\bar{g}_{a}gg_{a-1})$, then $(k,l,\bar{l}\,\bar{k}gkl,\bar{l}\,\bar{k}\,\bar{g}g_{a})$, then $(k,\bar{k}gk,l,\bar{l}\,\bar{k}\,\bar{g}kl)$, and finally $(g,k,l,\bar{l}\,\bar{k}\,\bar{g}\,g_{a})$ this becomes
$=\frac{\omega(k,\bar{k}gk,l)}{\omega(g,k,l)\,\omega(k,l,\bar{l}\,\bar{k}gkl)}|(g_{i})\rangle=\tau(\omega)_{g}(k,l)^{-1}\,|(g_{i})\rangle$
where $\tau(\omega)$ is the slant product of $\omega$, see Eq. (G.5). This shows that
Proposition J.1.
For any symmetry $\alpha$ the obstruction $\Lambda(\alpha)$ to covariant right restrictions is a function of its anomaly $\operatorname{\Omega}(\alpha)=[\omega]$ given by
$\Lambda(\alpha)=[\tau(\omega)]^{-1}$
where $\tau$ is the slant product, see Eq. (G.5).
Proof.
We verified the claimed equality for the example symmetries $\alpha_{\omega}$ with arbitrary anomaly $[\omega]$ constructed in Section D. If $\alpha$ is an arbitrary symmetry with anomaly $[\omega]$ then by Theorem B.1 it is stably equivalent to $\alpha_{\omega}$. By Proposition E.2 the obstruction $\Lambda$ is constant on stable equivalence classes. We conclude that $\Lambda(\alpha)=\Lambda(\alpha_{\omega})=[\tau(\omega)]^{-1}$. ∎
Appendix K Stable equivalence is necessary
We describe a ${\mathbb{Z}}_{2}$ symmetry with trivial anomaly which is not equivalent to a decoupled symmetry through conjugation by any FDQC. This shows that stacking with decoupled symmetries is really necessary to obtain the classification result of Theorem B.1.
Consider the spin-1/2 chain ${\mathcal{A}}$ with ${\mathcal{A}}_{j}\simeq\operatorname{\mathrm{End}}({\mathbb{C}}^{2})$ for all sites $j\in{\mathbb{Z}}$. For each site $j$ let $Z_{j}$ be the Pauli-Z operator at site $j$ and let $P^{\downarrow}_{j}=(\operatorname{\mathbbm{1}}-Z_{j})/2$. Write $CZ_{j,j^{\prime}}=\operatorname{\mathbbm{1}}-2P_{j}^{\downarrow}P_{j+1}^{\downarrow}$ for the controlled $Z$ gate acting on sites $j$ and $j^{\prime}$. Note that all these gates commute with each other.
For any finite interval $I=[a,b]$ define $U_{I}=\prod_{j=a}^{b-1}CZ_{j,j+1}$. Then $\alpha^{(-)}:=\lim_{a\uparrow\infty}\mathrm{Ad}[U_{[-a,a]}]$ defines the non-trivial component of a ${\mathbb{Z}}_{2}$-symmetry $\alpha$ with trivial anomaly.
Suppose it were possible to decouple $\alpha$ by a FDQC, then in particular there would exist a local unitary $V$ such that for all $a\in{\mathbb{N}}$ large enough we have $VU_{[-a,a]}V^{*}=U^{(L)}_{a}U^{(R)}_{a}$ with $U_{a}^{(L)}\in{\mathcal{A}}_{[-a,0]}$ and $U^{(R)}_{a}\in{\mathcal{A}}_{[1,a]}$.
To see that this is impossible, one first verifies by explicit computation that
$\operatorname{Tr}_{[0,1]}\left\{U_{[-a,a]}\right\}=2\times U_{[-a,-1]}\times CZ_{-1,2}\times U_{[3,a]}$
for all $a>3$, where $\operatorname{Tr}_{J}$ is the partial trace over ${\mathcal{A}}_{J}$ for any finite $J\subset{\mathbb{Z}}$. This is again a product of controlled Z’s so by an induction argument one obtains
$\operatorname{Tr}_{[-b,b+1]}\left\{U_{[-a,a]}\right\}=2^{b}\times U_{[-a,-(b+1)]}\times CZ_{-{b+1},b+2}\times U_{[b+2,a]}$ (K.1)
for all $a>b$. We can now take $a$ and $b$ large enough so that $V\in{\mathcal{A}}_{[-b,b]}$. By invariance of the partial trace under unitary conjugation we have $\operatorname{Tr}_{[-b,b]}\left\{VU_{[-a,a]}V^{*}\right\}=\operatorname{Tr}_{[-b,b+1]}\left\{U_{[-a,a]}\right\}$ is also given by Eq. (K.1). In contrast, if $VU_{[-a,a]}V^{*}=U^{(L)}_{a}U^{(R)}_{a}$ then $\operatorname{Tr}_{[-b,b]}\left\{VU_{[-a,a]}V^{*}\right\}=\operatorname{Tr}_{[-b,0]}\big{\{}U_{a}^{(L)}\big{\}}\times\operatorname{Tr}_{[1,b+1]}\big{\{}U_{a}^{(R)}\big{\}}$, which is incompatible with the form given by Eq. (K.1).